OpenStudy (anonymous):

ompare and Contrast: Below are two equations. Solve each equation and compare the two solutions. Choose the statement that is true about each solution. Equation #1 Equation #2 |x − 4| = 6 |3x + 12| = 18 Equation #1 and Equation #2 have the same number of solutions. Equation #1 has more solutions than equation #2. Equation #1 has fewer solutions than equation #2. None of the statements above describe the number of solutions to equations shown.

OpenStudy (anonymous):

x-4=+6 or -6 x-4=6 => x=2 x-4=-6 => x=-2 3x+12=18 or -18 3x+12=18 => x=2 3x+12=-18 => x=-10 Therefore, Equation #1 and Equation #2 have the same number of solutions.